OpenStudy (anonymous):

simply (4+i)(-8+10i) write answer in standard form.

OpenStudy (anonymous):

(4+i)(-8+10i) 4(8) + 4(10i) + i(-8) + i(10i) = 32 + 40i + -8i + 10i^2 = 10i^2 + 32i + 32 i hope that was helpful